OpenStudy (anonymous):

Chose the statement that is true? a) the value in column A is greater b) the value in column B is greater c) the two values are equal d) the relationship cannot be determined from the given information 9) Column A: the sum of the exterior angles of a 19-gon. Column B: the sum of the interior angles of a pentagon 10) Column A: the number of sides of a regular polygon with an exterior angle measuring 20 degrees Column B: the number of sides of a regular polygon whose sum of the interior angles measures 3240 degrees

OpenStudy (anonymous):

9) A)sum of exterior angle always 360degrees. B)sum of interior angle of a pentagon is: (5-2)*180= 540degrees. therefore statement (b) is true. 10) A)sum of exterior angle always 360degrees. therefore, number of sides= 360/20=18 B)let k be the number of sides. given sum of interior angles is 3240 degrees. so, (k-2)*180=3240. so, k= 20. therefore statement (b) is true